Suppose patient demand for blood tests at a local hospital to screen for various illnesses is given by Q-5,000-10P, where Q is the number of tests and P is the price of each test in dollars. It costs the hospital a constant $250 to run each test. Assume the market for blood tests is perfectly competitive, so the price the hospital charges for each test is equal to its marginal cost. In that case, the number of blood tests that will be run at the hospital is The deadweight loss that results from the hospital administering blood tests in this case is $ (enter only numbers in the blank, and please round to the nearest whole number if necessary) Question 17 4 pts Consider the hospital from the previous problem. Now suppose that all patients that visit the hospital have insurance that covers 80% of the cost of a blood test. In that case, the number of blood tests that will be run at the hospital is The deadweight loss that results from the hospital administering blood tests in this case is $ (enter only numbers in the blank, and please round to the nearest whole number if necessary) Question 18 2 pts Consider the hospital from the previous problems. Continue to assume that patients have insurance that covers 80% of the cost of a blood test. However, suppose that, because it helps to detect communicable diseases, each blood test that is administered has a positive externality that can be valued at $25. The deadweight loss resulting from the hospital administering blood tests in this case will be [size] in size relative to if there were no positive externality associated with the tests (enter just the word larger, smaller, or identical in the blank)
Answer #1

16. Under perfect competition;

Equilibrium is where P = MC

(5000 - Q)/10 = 250

5000 - Q = 2500

Q = 5000 - 2500

Q = 2500 units

No. of blood tests that will be run at the hospital is 2500.

DWL = Zero (under perfect competition there is no deadweight loss)